OpenStudy (anonymous):

What is the quotient (91y3 + 21y2 - 35y) ÷ 7y? @camerondoherty

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (mayankdevnani):

@SCARLET_SPIDER just divide them, \[\huge \bf \frac{91y^3+21y^2-35y}{7y}\]

OpenStudy (midhun.madhu1987):

divide each term by 7y separately

OpenStudy (anonymous):

kkk

OpenStudy (mayankdevnani):

you can also write it as :- \[\huge \bf \frac{91y^3}{7y}+\frac{21y^2}{7y}-\frac{35y}{7y}\]

OpenStudy (mayankdevnani):

as just midhun said.

OpenStudy (mayankdevnani):

can you now solve it ? @SCARLET_SPIDER

OpenStudy (anonymous):

Yes thanks @mayankdevnani

OpenStudy (mayankdevnani):

welcome :)